A car travels at 1/2 miles per hour for 8 minutes, then 1 mph for 4 minutes, then 2 mph for 2 minutes, then 4 miles per hour for 1 minute, and so forth: stopping after traveling 128 miles per hour for 1/32 minutes.How many miles does the car travel?
a) .56 b) .58 c) .6 d) .62 e) none of these

Answer by ankor@dixie-net.com(22740)

A car travels at 1/2 miles per hour for 8 minutes, then 1 mph for 4 minutes,
then 2 mph for 2 minutes, then 4 miles per hour for 1 minute, and so forth:
stopping after traveling 128 miles per hour for 1/32 minutes.
How many miles does the car travel?
:
Interesting that at each stage, the car travels the same distance; 1/15 of a mile
Converting minutes to hours:
.5(8/60) = (4/60) = (1/15)mi
1(4/60) = (1/15)mi
The last one
128(1/32)*(1/60) = (128/1920) = (1/15) mi
:
Total of all 9 stages
9(1/15) = .6 mi